OpenStudy (anonymous):

The function m = F(a) = a/8 gives the distance in inches on a map between two points that are actually a miles apart. 1. Whats if F(50)? What does F(50) represent?

OpenStudy (anonymous):

Is This Algebra 2?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

F(a) = a/8 F(50) = 50/8 ... replace each 'a' with 50 F(50) = ???

OpenStudy (anonymous):

6.25?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

`The function m = F(a) = a/8 gives the distance in inches on a map between two points that are actually a miles apart` so basically, whatever the value of 'a' is, the output F(a) tells us the distance on the paper map so in this case, a = 50 miles means F(a) = 6.25 inches 6.25 inches on the map corresponds to 50 miles in real life

OpenStudy (anonymous):

oh ok. and how about this?

OpenStudy (anonymous):

theres another part

OpenStudy (anonymous):

What are the units for the domain and range of the function? What is the domain and range for the function?

OpenStudy (anonymous):

Would the range be x> than the number?

jimthompson5910 (jim_thompson5910):

what are some possible values for 'a' ?

OpenStudy (anonymous):

50... 6.25...?

jimthompson5910 (jim_thompson5910):

a = 50 is one value, what's another?

OpenStudy (anonymous):

a = 8?

jimthompson5910 (jim_thompson5910):

sure, we could have some distance be 8 miles we could also have 2 miles, 1.5 miles, 99 miles basically any positive number or 0

OpenStudy (anonymous):

oh, really?

OpenStudy (anonymous):

so the domain is x>0

OpenStudy (anonymous):

Yeah He's right

jimthompson5910 (jim_thompson5910):

yeah \(\Large a \ge 0\)

OpenStudy (anonymous):

Oh yeah

OpenStudy (anonymous):

I forgot the equal to part... Do you think you could help me with two more parts?

jimthompson5910 (jim_thompson5910):

'a' is the input variable (not x this time) remember to underline the `>`

OpenStudy (anonymous):

Yeah I Believe The Domain Is X > 0

OpenStudy (anonymous):

Find the a in terms of m and lable this new function G. Tell what this function represents?

OpenStudy (anonymous):

would this be like a(m)?

jimthompson5910 (jim_thompson5910):

in terms of meters?

OpenStudy (anonymous):

i guess... It doesn't really say

jimthompson5910 (jim_thompson5910):

it seems like there's something missing but idk what

OpenStudy (anonymous):

It just says find a in terms of m and label this new function G. Tell what this function represents.

jimthompson5910 (jim_thompson5910):

maybe they just want you to convert to meters

jimthompson5910 (jim_thompson5910):

1 mile = 1609.34 meters (approximate)

OpenStudy (anonymous):

No, but I think it's just asking for a in terms of m. This is like functions and inverses

jimthompson5910 (jim_thompson5910):

1 mile = 1609.34 meters (approximate) a*1 mile = a*1609.34 meters .... Multiply both sides by a. a miles = 1609.34a meters

OpenStudy (anonymous):

I am so confused...

OpenStudy (anonymous):

I think (a)(m)

OpenStudy (anonymous):

Ok.. Let's try last part then. Determine if F and G are inverses

OpenStudy (anonymous):

But for that we need to find out what a is in terms of m.... :(

jimthompson5910 (jim_thompson5910):

yeah I'm not sure what they want for g(x)

OpenStudy (anonymous):

yeah... Ugh, this is so aggravating

OpenStudy (anonymous):

Let me see if I can figure it out, and If I still can't I'll call you back.. Ok..?

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

Thank you for your help so far though!!

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

You rang?

OpenStudy (anonymous):

Yes, thank you

OpenStudy (anonymous):

I just don't get this question... Do you mind leading me through it?

OpenStudy (mertsj):

Not at all. You found that the f(50)=6.25, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But from there, I just sorta lost it... I don't get the find a in terms of m thing

OpenStudy (mertsj):

Notice that the problem says that the function gives the distance in inches on a map.

OpenStudy (anonymous):

ok, I didn't think that was really important

OpenStudy (mertsj):

So 6.25 is the distance in inches on a map. Big deal you say.

OpenStudy (anonymous):

Ok. I get that much but then I don't

OpenStudy (mertsj):

And you have a point however, the second part of the question wants an explanation of the significance of 6.25 inches.

OpenStudy (anonymous):

Oh ok. Isn't the significance of it, the fact that you need it to be able to find G

OpenStudy (mertsj):

And the problem tells us that as well when it says that 6.25 inches is the distance on the map of two cities that are "a" miles apart and "a" in our problem is 50 because our function is f(a)=a/8 and we found f(50).

OpenStudy (anonymous):

Like, isn't the 6.25 the value of a? Or have I mistaken this entire thing....?

OpenStudy (anonymous):

So what can I do to find the domain and range?

OpenStudy (mertsj):

So we might say here is a map:

OpenStudy (anonymous):

Oh ok. So 50 is a? correct?